# Reinforcement Learning techniques for the flavor problem in particle physics

### v1

@@ 1 Introduction
The flavor problem in particle physics has been a central focus of study since the inception of the Standard Model (SM). Many attempts to explain fermion masses and mixing involve SM extensions based on horizontal symmetries, who are responsible of making the entries of the mass matrices related in a non trivial way instead of being free parameters. A common feature of all these models is the assignment of particles to irreducible representation of the flavor group (the case of non-abelian simmetries), the assignment of a $U(1)$ charge to particles (models á la Froggatt-Nielsen [1]) or a combination of both. The space of all possibilities is usually immense, spanning $\sim 10^{n}$ different models, with $n$ a large integer of $\mathcal{O}(10)$. It is clear that the success of an SM extension strongly relies on the ability to select a realization that fits the low energy data; while human searches have lead to a good success, it is not sure whether the discovered solutions are the minimal ones in terms of new degrees of freedom or the best ones according to other selection criteria like, for instance, the absence of huge hierarchies on the model parameters or a strong fine-tuning. The vastness of the parameters space in many fields of particle physics is one of the reason why, in very recent years, Machine Learning (ML) techniques started to be applied to several problems in Physics. Entering into the details of all attempts performed using ML techniques in Physics is a formidable task, as in many areas a long list of papers exist and revising them is well beyond the scope of this short review. In the restricted domain of theoretical physics, the subject is currently very active, covering a wide range of topics ranging (among others) from string theory [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] to QCD [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62], not to mention statistical mechanics [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90], thermodynamics [91, 92, 93, 94, 95], the foundation of quantum mechanics [96, 97, 98, 99, 100, 101, 102, 103, 104] and cosmology [105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121]. In this short review, we focus more strictly on ML in particle physics [122, 123, 124, 125, 126], with the intent of exploring Beyond the Standard Model theories [127, 128, 129] and, in particular, the (a few, so far) efforts that have been done to implement AI agents in the study of the flavor problem. The subject is relatively young so, at least in principle, there exists a large probability that more can be said and done in the near future. This review would help in setting the state of the art on the attempts accomplished using the Reinforcement Learning (RL) technique [130] to find promising quark [131] and lepton [132, 133] textures [^3] A study based on the Diffusion Model (DF) has been carried out for leptons [134]. addressing, at the same time, the quest for an axion candidate.
@@ 2 Reinforcement Learning
Reinforcement Learning [130] is a subfield of Machine Learning that focuses on how an agent learns to make optimal decisions by interacting with an environment to maximize a cumulative reward over time. Unlike supervised learning, where a model is trained on a dataset of labeled input-output pairs, or unsupervised learning, which seeks patterns in unlabeled data, RL operates in a trial-and-error framework. The agent learns by taking actions, observing their consequences, and receiving feedback in the form of rewards or penalties. This feedback loop allows the agent to gradually refine its decision-making strategy without being explicitly told what the “correct” action is.
RL draws inspiration from behavioral psychology, particularly the concept of operant conditioning, where behaviors are shaped by rewards and punishments. This paradigm is well-suited for problems where the optimal course of action is not immediately clear, and the agent must explore its environment to discover effective strategies. RL is often formalized using Markov Decision Processes (MDPs), which provide a mathematical framework to model sequential decision-making under uncertainty. In RL, the agent observes the current state of the environment, selects an action, and receives a reward along with the next state, iteratively learning a policy—a strategy that maps states to actions—to maximize the expected sum of future rewards.
@@ 2.1 Common features
  • Value Function ($V^{\pi}(s)$): a function that estimates the expected cumulative reward (also called the return) an agent can achieve starting from a given state $s$ and following the policy $\pi$. The value function quantifies the long-term desirability of a state under a specific policy. A related concept is the action-value function ($Q^{\pi}(s,a)$), which estimates the expected return for taking action $a$ in state $s$ and following policy $\pi$ afterward;
@@
The objective of RL is to find an optimal policy $\pi^{*}$ that maximizes the expected discounted return, defined as the expected sum of discounted rewards over an infinite horizon [130]:
@@
    Value-based methods: these methods estimate the value function ($V^{\pi}(s)$ or $Q^{\pi}(s,a)$) and derive a policy by selecting actions that maximize the estimated value. A classic example is Q-learning, which iteratively updates the action-value function to learn an optimal policy without explicitly modeling the environment’s transition dynamics [135]. At the heart of Q-learning is the Bellman equation used to update Q-values:
    $Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha\left[r_{t+1}+\gamma\max_{a}Q(s_{t+1},a)-Q(s_{t},a_{t})\right]\,$,
@@
    $Q(s,a)\leftarrow Q(s,a)+\alpha\left[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)\right]$
@@
    $Q^{*}(s,a)=\mathbb{E}\left[r_{t+1}+\gamma\max_{a^{\prime}}Q^{*}(s_{t+1},a^{\prime})\,\middle|\,s_{t}=s,a_{t}=a\right]\,.$
    In other words: the optimal value of taking action $a$ in state $s$ equals the expected value of the immediate reward $r_{t+1}$ plus the discounted ($\gamma$) value of the best possible future action in the next state $s_{t+1}$. The term $\max_{a^{\prime}}$” indicates that, in the future, the agent will always select the best action.
@@
    Policy-based methods: These methods directly parameterize and optimize the policy $\pi_{\theta}(a|s)$, where $\theta$ are tunable parameters of the policy (e.g., weights of a neural network) [136]. Instead of estimating a value function first, policy-based methods adjust $\theta$ to maximize the expected return:
    $J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t+1}\right].$
@@
    $\theta\leftarrow\theta+\alpha\sum_{t=0}^{\infty}\nabla_{\theta}\ln\pi_{\theta}(a_{t}|s_{t})G_{t},$
    where $G_{t}$ is the cumulative discounted reward from time $t$ defined in eq.(1). Policy-based methods naturally handle high-dimensional or continuous action spaces and can learn stochastic policies, which are useful in partially observable or highly uncertain environments.
There also exist hybrid approaches called actor-critic methods which combine value-based and policy-based ideas by learning both a value function (critic) and a policy (actor) to improve sample efficiency [137].
@@ 3.1 $U(1)$ Froggatt-Nielsen for quark observables
The first paper we want to discuss is Ref. [131], which explores the application of RL to model building, specifically focusing on Froggatt-Nielsen (FN) type models for quark masses and mixing. The primary goal is to investigate whether RL techniques can be used to train a neural network [138] to construct particle physics models with certain prescribed properties. More specifically, the authors aimed to develop a system where an RL agent could efficiently lead from random, typically physically unacceptable FN models, to phenomenologically viable ones that are consistent with observed quark masses and mixing. In addition, they also wanted to test if the trained networks could not only discover new models but also identify models already proposed in the literature.
@@
Table 1: Reference values of the quark masses (in GeV) and CKM parameters used in [131].
@@
The FN landscape consists of many possible charge assignments. For example, with one $U(1)$ and $q_{\rm max}=-q_{min}=9$, there are $19^{10}\sim 10^{13}$ models; for two $U(1)$ symmetries, $\sim 10^{26}$. Identifying viable models matching experimental data motivates the use of RL techniques. The setup involves a single neural network $\pi_{\theta}$, parameterized by weights $\theta$, which defines the policy $\pi$. The network takes states as input and outputs a probability distribution over possible actions. Exploration of the environment is driven by this policy, meaning actions are sampled according to $\pi_{\theta}$ at each step in an episode, defined as the sequence of states $s_{t}$ and actions $a_{t}$ for $t=0,1,2,$...:
@@
Table 2: Examples of models with highest intrinsic value for one (top) and two (bottom) $U(1)$ symmetries. Charges and $\mathcal{O}(1)$ coefficients of eq.(7) are given. Adapted from [131].
@@
Even though not strictly related to FN models, we want to mention here the efforts provided in [8] where machine learning techniques have been applied to construct particle physics models again in the Yukawa quark sector, aiming to satisfy two complementary criteria: truth and beauty. Truth is defined as the ability of a model to reproduce all experimental observables sensitive to its parameters. In the quark sector, this entails fitting quark masses and mixing parameters, including the CKM matrix elements and the Jarlskog invariant. Beauty [^6] The implementation of the beauty definitions in the context of machine learning approaches to BSM models in flavor physics might be useful to narrow down the viable models in order to search for the most theoretically appealing solutions., in the authors’ definition, refers to desirable theoretical attributes beyond mere data fitting. To quantify beauty in Yukawa matrices, three measures are tested, which are uniformity (that is all matrix elements have approximately equal magnitude), sparsity (where a significant number of elements vanish exactly or are negligibly small) and symmetry (in which off-diagonal elements are mirrored in magnitude across the main diagonal). Using pseudo-experiments, the optimization procedure efficiently navigates the huge parameter space, uncovering models that achieve both experimental accuracy and aesthetic elegance. This effort demonstrates how machine learning can be applied in particle physics, not only as an efficient method for exploring vast parameter spaces but also as a flexible tool for model selection (taking into account theoretical and phenomenological constraints), thus bridging numerical optimization and model-building intuition.
@@ 3.2 Extention of the $U(1)$ Froggatt-Nielsen to lepton observables
We now discuss the paper [132] which extends somehow the previous work to investigate the flavor structure not only for quarks but also for leptons using RL. Specifically, the authors apply a value-based RL algorithm to models with a $U(1)$ flavor symmetry based on the Froggatt-Nielsen (FN) mechanism. While, as said, for $U(1)$ charges $q_{i}$ in the range $9\leq q_{i}\leq 9$ to quarks alone yields approximately $10^{13}$ possible configurations, the inclusion of the lepton sector expands the parameter space to over $10^{24}$ possibilities. This combinatorial explosion renders brute-force searches infeasible and motivates the application of machine learning techniques for efficient model exploration. The system is provided with the following components:
@@
Table 3: Experimental values for the lepton sector obtained from global analysis of the data, where $\Delta m_{3l}^{2}\equiv\Delta m_{31}^{2}=m^{2}_{3}-m^{2}_{1}>0$ for NO and $\Delta m_{3l}^{2}\equiv\Delta m_{32}^{2}=m^{2}_{3}-m^{2}_{2}<0$ for IO. Taken from [132].
@@
Figure 1: Learning results for the lepton sector by RL having specified the neutrino mass ordering. Upper panel: loss function vs episode number. Lower panels: effective neutrino mass $m_{\beta\beta}$ (left plot) and sum of the neutrino masses (right plot) versus the atmospheric mixing angle $\theta_{23}$, for the 6 viable models found by the authors. See text for further details. Adapted from [132].
@@
Table 4: Highest intrinsic value scenario for the lepton sector with NO, where the neutrino mass ordering is specified in the network training. Adapted from [132].

### v2

@@ 1 Introduction
The flavor problem in particle physics has been a central focus of study since the inception of the Standard Model (SM). Many attempts to explain fermion masses and mixing involve SM extensions based on horizontal symmetries, who are responsible of making the entries of the mass matrices related in a non trivial way instead of being free parameters. A common feature of all these models is the assignment of particles to irreducible representation of the flavor group (the case of non-abelian symmetries), the assignment of a $U(1)$ charge to particles (models á la Froggatt-Nielsen [Froggatt:1978nt]) or a combination of both. The space of all possibilities is usually immense, spanning $\sim 10^{n}$ different models, with $n$ a large integer of $\mathcal{O}(10)$. It is clear that the success of an SM extension strongly relies on the ability to select a realization that fits the low energy data; while human searches have led to a good success, it is not sure whether the discovered solutions are the minimal ones in terms of new degrees of freedom or the best ones according to other selection criteria like, for instance, the absence of huge hierarchies on the model parameters or a strong fine-tuning. The vastness of the parameters space in many fields of particle physics is one of the reason why, in very recent years, Machine Learning (ML) techniques started to be applied to several problems in Physics. Entering into the details of all attempts performed using ML techniques in Physics is a formidable task, as in many areas a long list of papers exist and revising them is well beyond the scope of this short review. In the restricted domain of theoretical physics, the subject is currently very active, covering a wide range of topics ranging (among others) from string theory [Halverson:2019tkf, Mutter:2018sra, He:2017aed, Krefl:2017yox, Ruehle:2017mzq, Abel:2014xta, Matchev:2024ash, Larfors:2020ugo, Krippendorf:2021uxu, Constantin:2021for, Cole:2021nnt, Seong:2024wkt, Harvey:2024lwl, Lanza:2024mqp, Choi:2023rqg, Hirst:2022jfq, Cui:2022cxe, Bao:2022rup, Berman:2021mcw, Gao:2021xbs, Carifio:2017nyb, Liu:2017dzi, Wang:2018rkk, Altman:2018zlc, Jinno:2018dek, Bull:2018uow, Jejjala:2019kio, Brodie:2019dfx, Suresh:2024dsg, Bies:2020gvf, Ruehle:2020jrk, Anderson:2020hux, Cipriani:2025ini] to QCD [Shanahan:2018vcv, BarreraCabodevila:2025iwr, BarreraCabodevila:2025ogv, Favoni:2020reg, Favoni:2025ghc, Holland:2025fsa, Zhu:2025pmw, Aarts:2025gyp, Tomiya:2025quf, Christ:2024ydh, Gao:2024nzg, Apte:2024vwn, Cruz:2024grk, Abbott:2024kfc, Chatterjee:2024pbp, Holland:2024muu, Gao:2023quv, Gao:2023uel, Ermann:2023unw, Zhou:2023pti, He:2023zin, Lehner:2023bba, Shanahan:2022ifi, Urban:2022vjz, Boyda:2022nmh, Gerasimeniuk:2021cql], not to mention statistical mechanics [Carleo:2016svm, Wetzel:2017ooo, Schlomer:2025txn, Du:2025qph, Wetzel:2025uhj, Zhang:2024mqd, Suresh:2024nek, Cybinski:2024wgx, Sadoune:2024hxl, Kairon:2023adn, Jianmin:2023iyz, Baul:2023qmu, Naravane:2023hbr, Arnold:2023tsr, Ghosh:2023hsg, Schlomer:2023bbd, Chung:2023uuj, Johnston:2022cgm, Patel:2022clf, Sancho-Lorente:2021yko, Chung:2021eoy, Kim:2021ulz, Cole:2020hjx, Dawid:2020itb, Alexandrou:2019hgt, Dong:2019ypp, Giannetti:2018vif, Nomura:2017tgw], thermodynamics [Torlai:2016szv, ShibaFunai:2018aaw, Seif:2019wdb, Zhang:2024wgr, Gu:2022gcn], the foundation of quantum mechanics [Deng:2017uik, Chen:2022ytr, Bharti:2020gua, Gonzalez:2020frf, Alves:2024wwu, Deng:2017hws, Bharti:2019xdc, Canabarro:2018xwk, Krivachy:2020ggy] and cosmology [Rudelius:2018yqi, Jejjala:2020wcc, Saxena:2024rhu, Ocampo:2024fvx, Savchenko:2024zyt, Decant:2024bpg, Savchenko:2025jzs, Hai-LongHuang:2025vfs, Eckner:2025waa, Villaescusa-Navarro:2021pkb, Hassan:2021ymv, Modi:2021acq, Modi:2020dyb, Dai:2020ekz, Taylor:2019mgj, Schmittfull:2017uhh, Flauger:2022hie]. In this short review, we focus more strictly on ML in particle physics [Kawai:2024pws, Schwartz:2021ftp, Abdelhaq:2025nch, Bakshi:2025fgx, Boto:2025ovp], with the intent of exploring Beyond the Standard Model theories [Wojcik:2024lfy, Wojcik:2024azu, Chekanov:2025pmb] and, in particular, the (a few, so far) efforts that have been done to implement AI agents in the study of the flavor problem. The subject is relatively young so, at least in principle, there exists a large probability that more can be said and done in the near future. This review would help in setting the state of the art on the attempts accomplished using the Reinforcement Learning (RL) technique [Sutton] to find promising quark [Harvey:2021oue] and lepton [Nishimura:2020nre, Baretz:2025zsv] textures [^3] A study based on the Diffusion Model (DF) has been carried out for leptons [Nishimura:2025knz]. addressing, at the same time, the quest for an axion candidate. More generally, the same RL-based strategy can be straightforwardly adapted to other areas of particle physics, such as to problems in theoretical physics where model building amounts to exploring large, discrete or continuous spaces under theoretical and phenomenological constraints. Among others, we mention string theory [Abel:2021rrj], particle jets physics [Carrazza:2019efs], lattice QCD [Alvestad:2023jgl] and quantum field theory [Windisch:2019byg, Zeng:2025xbh].
@@ 2 Reinforcement Learning
Reinforcement Learning [Sutton] is a subfield of Machine Learning that focuses on how an agent learns to make optimal decisions by interacting with an environment to maximize a cumulative reward over time. Unlike supervised learning, where a model is trained on a dataset of labeled input-output pairs, or unsupervised learning, which seeks patterns in unlabeled data, RL operates in a trial-and-error framework, like a human strategist attempting to master a complex task. The agent learns by taking actions, observing their consequences, and receiving feedback in the form of rewards or penalties. This feedback loop allows the agent to gradually refine its decision-making strategy without being explicitly told what the “correct” action is. This continuous refinement process, based on performance feedback rather than explicit instruction, mirrors how humans develop expertise in complex domains. This paradigm is well-suited for problems where the optimal course of action is not immediately clear, and the agent must explore its environment to discover effective strategies. RL is often formalized using Markov Decision Processes (MDPs), which provide a mathematical framework to model sequential decision-making under uncertainty. In RL, the agent observes the current state of the environment, selects an action, and receives a reward along with the next state, iteratively learning a policy—a strategy that maps states to actions—to maximize the expected sum of future rewards.
@@ 2.1 Common features
  • Value Function ($V^{\pi}(s)$): a function that estimates the expected cumulative reward (also called the return) an agent can achieve starting from a given state $s$ and following the policy $\pi$. The value function quantifies the long-term desirability of a state under a specific policy. A related concept is the action-value function ($Q^{\pi}(s,a)$), which estimates the expected return for taking action $a$ in state $s$ and following policy $\pi$ afterward. Notice that the reward $r$ and the value function are fundamentally related: the reward provides the immediate feedback signal, while the value function is the estimation of the long-term benefit derived from those rewards.
@@
The objective of RL is to find an optimal policy $\pi^{*}$ that maximizes the expected discounted return, defined as the expected sum of discounted rewards over an infinite horizon [Sutton]:
@@
    Value-based methods: these methods estimate the value function ($V^{\pi}(s)$ or $Q^{\pi}(s,a)$) and derive a policy by selecting actions that maximize the estimated value. A classic example is Q-learning, which iteratively updates the action-value function to learn an optimal policy without explicitly modeling the environment’s transition dynamics [Wang:2021yix]. At the heart of Q-learning is the Bellman equation used to update Q-values:
    $Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha\left[r_{t+1}+\gamma\max_{a}Q(s_{t+1},a)-Q(s_{t},a_{t})\right]\,$, (2)
@@
    $Q(s,a)\leftarrow Q(s,a)+\alpha\left[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)\right]$ (3)
@@
    $Q^{*}(s,a)=\mathbb{E}\left[r_{t+1}+\gamma\max_{a^{\prime}}Q^{*}(s_{t+1},a^{\prime})\,\middle|\,s_{t}=s,a_{t}=a\right]\,$, (4)
    where the symbol $\mathbb{E}$ defines the expected value. In other words: the optimal value of taking action $a$ in state $s$ equals the expected value of the immediate reward $r_{t+1}$ plus the discounted ($\gamma$) value of the best possible future action in the next state $s_{t+1}$. The term $\max_{a^{\prime}}$” indicates that, in the future, the agent will always select the best action.
@@
    Policy-based methods: These methods directly parameterize and optimize the policy $\pi_{\theta}(a|s)$, where $\theta$ are tunable parameters of the policy (e.g., weights of a neural network) [Sequeira:2024htb]. Instead of estimating a value function first, policy-based methods adjust $\theta$ to maximize the expected return:
    $J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t+1}\right].$ (6)
@@
    $\theta\leftarrow\theta+\alpha\sum_{t=0}^{\infty}\nabla_{\theta}\ln\pi_{\theta}(a_{t}|s_{t})G_{t},$ (7)
    where $G_{t}$ is the cumulative discounted reward from time $t$ defined in eq.(1). REINFORCE is suitable for RL because it estimates the policy gradient from sampled trajectories without requiring a model of the environment. However, its gradient estimates have high variance, often necessitating variance-reduction techniques for stable and efficient learning.
    Policy-based methods naturally handle high-dimensional or continuous action spaces and can learn stochastic policies, which are useful in partially observable or highly uncertain environments.
There also exist hybrid approaches called actor-critic methods which combine value-based and policy-based ideas by learning both a value function (critic) and a policy (actor) to improve sample efficiency [Kolle:2024lpx].
@@ 3.1 $U(1)$ Froggatt-Nielsen for quark observables
The first paper we want to discuss is Ref. [Harvey:2021oue], which explores the application of RL to model building, specifically focusing on Froggatt-Nielsen (FN) type models for quark masses and mixing. The primary goal is to investigate whether RL techniques can be used to train a neural network [Nielsen] to construct particle physics models with certain prescribed properties. More specifically, the authors aimed to develop a system where an RL agent could efficiently lead from random, typically physically unacceptable FN models, to phenomenologically viable ones that are consistent with observed quark masses and mixing. In addition, they also wanted to test if the trained networks could not only discover new models but also identify models already proposed in the literature.
@@
Table 1: Reference values of the quark masses (in GeV) and CKM parameters used in [Harvey:2021oue].
@@
The FN landscape consists of many possible charge assignments. For example, with one $U(1)$ and $q_{\rm max}=-q_{min}=9$, there are $19^{10}\sim 10^{13}$ models; for two $U(1)$ symmetries, $\sim 10^{26}$. Identifying viable models matching experimental data motivates the use of RL techniques. The setup involves a single neural network $\pi_{\theta}$, parameterized by weights $\theta$, which defines the policy $\pi$. The network takes states as input and outputs a probability distribution over possible actions. Exploration of the environment is driven by this policy, meaning actions are sampled according to $\pi_{\theta}$ at each step in an episode, defined as the sequence of states $s_{t}$ and actions $a_{t}$ for $t=0,1,2,$... (the maximum $t$ representing the length of the episode):
@@
Table 2: Examples of models with highest intrinsic value for one (top) and two (bottom) $U(1)$ symmetries. Charges and $\mathcal{O}(1)$ coefficients of eq.(13) are given. Adapted from [Harvey:2021oue].
@@
Even though not strictly related to FN models, we want to mention here the efforts provided in [Matchev:2024ash] where machine learning techniques have been applied to construct particle physics models again in the Yukawa quark sector, aiming to satisfy two complementary criteria: truth and beauty. Truth is defined as the ability of a model to reproduce all experimental observables sensitive to its parameters. In the quark sector, this entails fitting quark masses and mixing parameters, including the CKM matrix elements and the Jarlskog invariant. Beauty [^6] The implementation of the beauty definitions in the context of machine learning approaches to BSM models in flavor physics might be useful to narrow down the viable models in order to search for the most theoretically appealing solutions., in the authors’ definition, refers to desirable theoretical attributes beyond mere data fitting. To quantify beauty in Yukawa matrices, three measures are tested, which are uniformity (that is all matrix elements have approximately equal magnitude), sparsity (where a significant number of elements vanish exactly or are negligibly small) and symmetry (in which off-diagonal elements are mirrored in magnitude across the main diagonal). Using pseudo-experiments, the optimization procedure efficiently navigates the huge parameter space, uncovering models that achieve both experimental accuracy and aesthetic elegance. This effort demonstrates how machine learning can be applied in particle physics, not only as an efficient method for exploring vast parameter spaces but also as a flexible tool for model selection (taking into account theoretical and phenomenological constraints), thus bridging numerical optimization and model-building intuition.
@@ 3.2 Extention of the $U(1)$ Froggatt-Nielsen to lepton observables
We now discuss the paper [Nishimura:2020nre] which extends somehow the previous work to investigate the flavor structure not only for quarks but also for leptons using RL. Specifically, the authors apply a value-based RL algorithm to models with a $U(1)$ flavor symmetry based on the Froggatt-Nielsen (FN) mechanism. While, as said, for $U(1)$ charges $q_{i}$ in the range $9\leq q_{i}\leq 9$ to quarks alone yields approximately $10^{13}$ possible configurations, the inclusion of the lepton sector expands the parameter space to over $10^{24}$ possibilities. This combinatorial explosion renders brute-force searches infeasible and motivates the application of machine learning techniques for efficient model exploration. The system is provided with the following components:
@@
Table 3: Experimental values for the lepton sector obtained from global analysis of the data, where $\Delta m_{3l}^{2}\equiv\Delta m_{31}^{2}=m^{2}_{3}-m^{2}_{1}>0$ for NO and $\Delta m_{3l}^{2}\equiv\Delta m_{32}^{2}=m^{2}_{3}-m^{2}_{2}<0$ for IO. Taken from [Nishimura:2020nre].
@@
Figure 1: Learning results for the lepton sector by RL having specified the neutrino mass ordering. Upper panel: loss function vs episode number. Lower panels: effective neutrino mass $m_{\beta\beta}$ (left plot) and sum of the neutrino masses (right plot) versus the atmospheric mixing angle $\theta_{23}$, for the 6 viable models found by the authors. See text for further details. Adapted from [Nishimura:2020nre].
@@
Table 4: Highest intrinsic value scenario for the lepton sector with NO, where the neutrino mass ordering is specified in the network training. Adapted from [Nishimura:2020nre].